OpenStudy (anonymous):

Eight white balls and twenty black balls are in a bag. What is the probability of drawing a black ball in one draw? 1 2/5 5/7

OpenStudy (anonymous):

ok so, to solve this problem first we will need to find the ratio between the white balls and the black balls.

OpenStudy (anonymous):

there are 8 white and 20 black tell me the ratio of the white balls to the black balls

OpenStudy (anonymous):

8:20

OpenStudy (anonymous):

correct! now we need to simplify the ratio... hold on...

OpenStudy (anonymous):

ok, now to simplify. what to factors equal 8?

OpenStudy (anonymous):

here... im sure your in a hury so I will tell you 8= 2*4 20=4*5

OpenStudy (anonymous):

then we cross out the "like factors" which is basically crossing out the two twos P.S 20=2*10

OpenStudy (anonymous):

then you have 4/5 as your answer

OpenStudy (anonymous):

got that?

OpenStudy (anonymous):

oh wait! its talking about black balls specifically I read this wrong it is actually 2/7

OpenStudy (anonymous):

im sorry im back